Operations Report – March 8th

Crew 206 Operations Report 08 – MAR – 2019

SOL: 12

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): 12,3 hours

Solar— SOC % – Turned off at 79%, Turned on at 80% (18h30)

SOC tracking :
7h45: 78%
9h30: 74%
11h30: 98%
13h40: 92%
16h40: 86%
17h30: 82%
18h30: 80%

Diesel Reading – 65%
Propane Reading – 78%

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 400 gallons

Water differential (static tank) – 9 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 232 gallons

Static to Loft Pump used – Yes

Water Meter: 0141699,3 gallons

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Breaking system tested.

Hours: 53,1 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

Curiosity rover used: Not Used

Hours: 78.0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 78,9 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Opportunity’s brake fluid is empty, it looks like the
pipeline has a leak. Atila is working on it, more informations will follow.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Used for Staff.

Staff Crew Car used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: Still investigating. Nothing more to report today.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: None of the tarps seem to have suffered the
high winds today, everything else is doing fine too.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.